# Generating Diverse and High-Quality Texts by Minimum Bayes Risk Decoding

### v1

@@ 4.2 Image Captioning using BLIP-2
(a) P-SentBERT $\downarrow$ (MS COCO)
(b) P-SentBERT $\downarrow$ (SQuADv2)
(c) P-SentBERT $\downarrow$ (CommonGen)
Figure 5: Evaluation of semantic textual similarity using pairwise Sentence BERT as a function of the mean quality score (BLEU and METEOR). The number of outputs $k$ is $4$.
(a) P-SentBERT $\downarrow$ (MS COCO)
(b) P-SentBERT $\downarrow$ (SQuADv2)
(c) P-SentBERT $\downarrow$ (CommonGen)
Figure 6: Evaluation of semantic textual similarity using pairwise Sentence BERT as a function of the Oracle quality score (max BLEU and max METEOR). The number of outputs $k$ is $4$.
We evaluate the performance of the proposed methods on image captioning using MS COCO dataset Lin et al. (2014). We use BLIP-2 Li et al. (2023) with Flan T5-xl Chung et al. (2022) fine-tuned for MS COCO loaded in 8-bit. [^4] https://huggingface.co/Salesforce/blip2-flan-t5-xl-coco The number of outputs $k$ is set to $4,8$, and $12$. We evaluate the performance of epsilon sampling with $\epsilon=0.02$. The diversity penalty for DBS is set to $0.0,0.2,0.5,1.0,2.0$, and $5.0$. For DMBR, we generate $64$ samples using epsilon sampling with $\epsilon=0.02$. The results on $k=4$ are shown in Figure 4. DMBR achieves lower pairwise-BLEU and higher distinct-n than DBS and epsilon sampling. See Appendix C for the other experimental results.
Semantic Textual Similarity
While the machine translation task requires generating semantically the same sentence, the image captioning task allows more diversity in the semantics of the output. In addition to the lexical diversity metrics, we evaluate the semantic diversity of the generated outputs using sentence BERT Reimers and Gurevych (2019). We compute the embedding of each output using the sentence BERT and compute the cosine similarity of each pair of outputs. The cosine similarity over a pair of sentences is shown to have a high correlation to human preference Tevet and Berant (2021). We evaluate pairwise-sentBERT, the average cosine similarity over a set of pairs of outputs. We use all-mpnet-base-v2 model. [^5] https://huggingface.co/sentence-transformers/all-mpnet-base-v2 The model is based on MPNet Song et al. (2020) and has shown to be effective for a variety of sentence embedding tasks. [^6] https://www.sbert.net/docs/pretrained_models.html The result in Figure 4(a) shows that DMBR achieves better (lower) pairwise-sentBERT than DBS and epsilon sampling.
4.3 Question Generation using Large Language Model
The goal of question generation is to generate a question on a topic in natural language from a paragraph of text Mulla and Gharpure (2023). We use a Stanford Question Answering Dataset (SQuAD2.0) to evaluate the decoding algorithms for question generation Rajpurkar et al. (2016, 2018). SQuAD2.0 is a reading comprehension dataset consisting of questions and answers on Wikipedia articles. We use Zephyr-7B $\beta$ with prompting as a text generation model. [^7] https://huggingface.co/HuggingFaceH4/zephyr-7b-beta We use the following prompt:
@@ 4.3 Question Generation using Large Language Model
We generate $k=4,8,12$ outputs. For DBS, we set the diversity penalty to $0.0,0.5,1.0,2.0,$ and $5.0$. For MBR we generate $128$ samples with epsilon sampling with $\epsilon=0.01$. The diversity penalty $\lambda$ is set to $0.0,0.1,0.3,0.5,1.0,$ and $2.0$. The mean METEOR and the diversity metrics are shown in Figure 4. Compared with the same METEOR score, DMBR has better distinct-2 and 3 than DBS. The pairwise-BLEU and pairwise-sentBERT (Figure 5) of DMBR are slightly worse than DBS and epsilon sampling compared with the same mean METEOR score.
4.4 Generative Common Sense Reasoning using Large Language Model
@@
The other experimental setting is the same as in Section 4.3. The mean METEOR and the diversity metrics are shown in Figure 4. DMBR has better distinct-2 and 3 than DBS and epsilon sampling but slightly worse pairwise-BLEU.
Coverage
The task encourages the generated sentence to contain as many input concepts as possible, so we also consider the coverage of the input concepts. The definition of coverage is the number of captured concepts divided by the number of input concepts. To compute the captured concepts, we use a Porter stemmer Porter (1980) to extract the stems of the input concepts and the generated sequences and compute the number of overlaps. [^8] https://www.nltk.org/_modules/nltk/stem/porter.html Overall, the coverage is low ($<18\%$) for all the decoding algorithms. For $k=4$, the average coverage is $15.56,15.42,16.03,15.12,$ and $15.93$ for beam search, DBS-1.0, MBR, DMBR-1.0, and KMBR, respectively. We speculate this is because the text generation model is not trained to follow the instructions on the lexical constraints without few-shot prompting.
@@ 4.5 Text Summarization
We use the Extreme Summarization (XSum) dataset as a benchmark for an abstractive single-document summarization Narayan et al. (2018). We use a BART model pretrained on XSum dataset Lewis et al. (2020). [^9] https://huggingface.co/facebook/bart-large-xsum We evaluate the quality of the outputs by ROUGE-L using HuggingFace’s evaluate library Lin (2004). We generate $k=4$ outputs per each input document. For DBS, we set the diversity penalty to $0.0,0.5,1.0,2.0,5.0$. For MBR we generate 64 samples with epsilon sampling with $\epsilon=0.02$. The diversity penalty $\lambda$ for DMBR is set to $0.0,0.1,0.3,0.5,1.0,2.0$. The results are shown in Figure 4. DMBR achieves better diversity than DBS measured by P-BLEU and distinct-n.
5 Related Work
Several work have investigated strategies for increasing the diversity of text generation by modifying the decoding algorithm. Many of the studies have proposed modifications to beam search to promote diverse set of beams Vijayakumar et al. (2016); Li and Jurafsky (2016); Kulikov et al. (2019); Kool et al. (2019); Tam (2020); Meister et al. (2021b, a). Beam search-based diverse decoding algorithms are often deployed for directed text generation tasks such as image captioning. Some other work has proposed extension to random sampling algorithm to increase quality of the output at a cost of diversity Fan et al. (2018); Tu et al. (2019); Holtzman et al. (2020); Zhang et al. (2021); Hewitt et al. (2022); Vilnis et al. (2023).
Our work is most closely related to the Post-Decoding Clustering (PDC) approach Kriz et al. (2019); Ippolito et al. (2019). Similarly to DMBR, PDC generates a set of outputs and then chooses which ones to select as the final set of outputs. Kriz et al. (2019) proposes to cluster similar sentences by K-means and select outputs closest to the centroid of each cluster as the final outputs. Ippolito et al. (2019) instead select outputs with the highest log-likelihood from each cluster.
MBR decoding has been studied in many NLP tasks including parsing Goodman (1996), speech recognition Goel and Byrne (2000), bilingual word alignment Kumar and Byrne (2002), and machine translation Kumar and Byrne (2004). It has recently gained attention in machine translation as a decision rule to overcome some of the biases of MAP decoding in neural machine translation Eikema and Aziz (2020); Müller and Sennrich (2021); Eikema and Aziz (2022). Although it is shown to generate high-quality texts in directed text generation tasks such as machine translation, using MBR to generate a diverse set of outputs is yet to be investigated.
@@ 6 Conclusions
We study the problem of generating a set of texts with high quality and diversity. Our approach is to promote diversity to MBR which is shown to generate high quality texts so that it can generate high quality and diversity outputs. We extend MBR and propose DiverseMBR and KMBR that seek to optimize both the diversity and the quality when selecting a set of outputs. Because both algorithms takes exponential time to compute exactly, we devise approximate algorithms to make it feasible. We evaluate DMBR and KMBR on machine translation, image captioning, question generation, generative common sense reasoning, and text summarization tasks and show that overall they achieve better quality-diversity trade-off than DBS. We also observe that both methods, especially KMBR, achieve higher maximum BLEU score than MBR, analoguous to DBS achieving higher BLEU score than beam search.
@@ 7 Limitations
Our experiments are focused on directed text generation tasks. The evaluation of the methods in open-ended text generation tasks such as story generation is future work.
We rely on automatic evaluations to evaluate the quality and the diversity of the generated texts. Human evaluation is desirable, especially for evaluating diversity. Although the automatic evaluation metrics for diversity used in this paper (e.g. pairwise-BLEU, distinct-n, sentence-BERT) are shown to correlate with human evaluation, there is still a clear gap between automatic metrics and humans Tevet and Berant (2021).
DMBR and KMBR are much slower than DBS as they require the computation of the MBR objective which needs a computation of the utility function for $N^{2}$ times. Although recent work has shown that the computation of the MBR objective can be significantly reduced Cheng and Vlachos (2023); Jinnai and Ariu (2024), it is not directly applicable to DMBR and KMBR. How to reduce the inference time of these algorithms is future work.
@@
We consider the Monte Carlo estimate (Eq. 7) as the target quality objective for simplicity. Using other quality objective functions such as model-based estimate Jinnai et al. (2023) is future work.
@@ Acknowledgements
We thank Naoto Ohsaka for providing insights on approximation algorithms for a non-monotonic submodular function maximization problem.
@@ Appendix A Proof of Submodularity
Appendix B Examples of Generations
We show outputs generated by the decoding algorithms. Table 1, 2, 3, 4, 5, and 6 are the examples of the generations with various sampling algorithms for each domain evaluated in Section 4. The examples show the generations of the first input source of each dataset.
@@ Appendix B Examples of Generations
Appendix C Additional Results
The evaluation of the diversity as a function of minimum BLEU over the outputs on WMT’19 datasets is present in 7. DMBR achieves a better trade-off than DBS and sampling algorithms with the same minimum BLEU score.
The Oracle (maximum) and the minimum quality scores on MS COCO, SQuADv2, CommonGen, and XSum with an output size of 4 are present in Figure 8 and 9. We observe similar trends as in machine translation tasks.
Figure 10 shows the pairwise sentence BERT as a function of the minimum BLEU and METEOR scores on MS COCO, SQuAD, and CommonGen. The result indicates that DMBR and DBS are successfully generating semantically diverse outputs.
Table 7, 8, 9, 10, 11, 12, and 13 show the summary of the experimental results described in Section 4.
@@ Appendix C Additional Results
Figure 7: Min BLEU, pairwise-BLEU, distinct-n on WMT’19 De-En and Ru-En. The size of the output $k$ is set to 4.
@@
Figure 9: Evaluation of the pairwise-BLEU and distinct-2, 3 as a function of min BLEU (MS COCO), METEOR (SQuAD v2, CommonGen), and ROUGE-L (XSum). The size of the output $k$ is set to 4.
@@
Figure 10: Evaluation of semantic textual similarity using pairwise Sentence BERT as a function of the Oracle quality score (min BLEU and min METEOR). The number of outputs $k$ is $4$.

### v2

@@ 4.2 Image Captioning using BLIP-2
We evaluate the performance of the proposed methods on image captioning using MS COCO dataset Lin et al. (2014). We use BLIP-2 Li et al. (2023) with Flan T5-xl Chung et al. (2022) fine-tuned for MS COCO. We load the model in 8-bit to reduce the VRAM consumption. The number of outputs $k$ is set to $\{4,8,12\}$. We evaluate the performance of epsilon sampling with $\epsilon\in\{0.01,0.02,0.04\}$. The diversity penalty for DBS is set to $\{0.0,0.2,0.5,1.0,2.0,5.0\}$. For DMBR, we generate $N=64$ samples using epsilon sampling with $\epsilon=0.02$. The results on $k=4$ are shown in Figure 6 (6(a), 6(d), 6(g)). DMBR achieves lower P-BLEU and higher distinct-2 than DBS and epsilon sampling.
Evaluation of Semantic Diversity.
While the machine translation task requires generating a text semantically the same to the input text, the image captioning task allows more diversity in the contents of the output Wang and Chan (2019). To evaluate the semantic diversity beyond the surface diversity, we employ sentence BERT Reimers and Gurevych (2019). We compute the embedding of each output using the sentence BERT and compute the cosine similarity of each pair of outputs. The cosine similarity over a pair of sentences is shown to have a high correlation to human preference Tevet and Berant (2021). We evaluate the pairwise sentence BERT (P-SentBERT), the average cosine similarity of the sentence embeddings over a set of pairs of outputs. We use all-mpnet-base-v2 model. The model is based on MPNet Song et al. (2020) and has shown to be effective for a variety of sentence embedding tasks. The result in Figure 6(g) shows that DMBR achieves better (lower) P-SentBERT than DBS and epsilon sampling.
4.3 Question Generation using Language Model
The goal of question generation is to generate a question on a topic in natural language from a paragraph of text Mulla and Gharpure (2023). We use a Stanford Question Answering Dataset (SQuADv2) to evaluate the decoding algorithms for question generation Rajpurkar et al. (2016, 2018). SQuADv2 is a reading comprehension dataset consisting of questions and answers on Wikipedia articles. We use a language model Zephyr-7B $\beta$ Tunstall et al. (2023) with prompting as a text generation model. We use the following prompt:
@@ 4.3 Question Generation using Language Model
We generate $k\in\{4,8,12\}$ outputs. For DBS, we set the diversity penalty to $\{0.0,0.5,1.0,2.0,5.0\}$. For MBR we generate $128$ samples with epsilon sampling with $\epsilon=0.01$. The diversity penalty $\lambda$ for DMBR is set to $\{0.0,0.1,0.3,0.5,1.0,2.0\}$. The mean METEOR and the diversity metrics are shown in Figure 6 (6(b), 6(e), 6(h)). Compared with the same METEOR score, DMBR has better distinct-2 and P-SentBERT than DBS. The P-BLEU and P-sentBERT of DMBR are slightly worse than DBS and epsilon sampling compared with the same mean METEOR score.
We speculate that DMBR underperforms DBS for SQuADv2 and CommonGen (Section 4.4) because DBS is more likely to generate a set of sequences of varying lengths. Table 1 shows the average standard deviation of the sequences generated by DMBR and DBS with varying diversity penalties. DBS tends to generate a set of diverse lengths of sentences than DMBR in these domains. Because P-BLEU is sensitive to the difference in length whereas distinct-n and P-SentBERT are less sensitive, P-BLEU is a favorable metric for DBS. As such, DMBR underperforms DBS for SQuADv2 and CommonGen in P-BLEU but not in distinct-n and P-SentBERT. It implies that if a practitioner benefits from varying sequence lengths, DBS may be preferred over DMBR, and if not, DMBR may be preferred over DBS.
DMBR
$\lambda$ 0.1 0.5 1.0 2.0
SQuADv2 3.73 4.29 4.30 5.38
CommonGen 3.36 3.69 4.41 6.86
DBS
$\lambda$ 0.5 1.0 2.0 5.0
SQuADv2 2.70 4.72 6.37 5.96
CommonGen 1.92 5.46 7.50 7.45
Table 1: The standard deviation of the sequence lengths averaged over the inputs using DMBR and DBS. Note that the diversity strength $\lambda$ of the two algorithms are used differently.
4.4 Generative Common Sense Reasoning using Language Model
@@
The other experimental setting is the same as in Section 4.3. The mean METEOR and the diversity metrics are shown in Figure 6 (6(c), 6(f), 6(i)). DMBR has better distinct-2 than DBS and epsilon sampling but slightly worse P-BLEU. We observe that the coverage of the input concepts is low ($<18\%$) in all settings in our experiments. See Appendix B for the analysis of the coverage.
(a) P-BLEU $\downarrow$ (XSum)
(b) distinct-2 $\uparrow$ (XSum)
Figure 7: Evaluation of the P-BLEU and distinct-2 as a function of mean ROUGE-L. The number of the outputs $k$ is $4$.
@@ 4.5 Text Summarization
We use the XSum dataset as a benchmark for an abstractive single-document summarization Narayan et al. (2018). We use a BART model pretrained on XSum dataset Lewis et al. (2020). We evaluate the quality of the outputs by ROUGE-L using HuggingFace’s evaluate library Lin (2004). We generate $k=4$ outputs per each input document. For DBS, we set the diversity penalty to $\{0.0,0.5,1.0,2.0,5.0\}$. For MBR we generate 64 samples with epsilon sampling with $\epsilon=0.02$. The diversity penalty $\lambda$ for DMBR is set to $\{0.0,0.1,0.3,0.5,1.0,2.0\}$. The results are shown in Figure 7. DMBR achieves better diversity than DBS measured by P-BLEU and distinct-n.
@@ 5 Conclusions
We study the problem of generating a set of texts with high quality and diversity. Our approach is to promote diversity to MBR which is shown to generate high quality texts so that it can generate high quality and diverse outputs. We extend MBR and propose DMBR and KMBR that seek to optimize both the diversity and the quality when selecting a set of outputs. Because both algorithms are too expensive to compute exactly, we devise approximate algorithms to make it feasible. We evaluate DMBR and KMBR on machine translation, image captioning, question generation, generative common sense reasoning, and text summarization tasks and show that overall they achieve better quality-diversity trade-off than DBS. We also observe that both methods, especially KMBR, achieve a higher max BLEU score than MBR, analogous to DBS achieving a higher BLEU score than beam search.
@@ 6 Limitations
Our experiments are focused on directed text generation tasks. Open-ended and directed text generation tasks are different tasks for text generation algorithms. This distinction separates what kind of text generation algorithms are suitable for each task. While beam search variants tend to perform better in directed text generation tasks, stochastic decoding algorithms tend to do better in open-ended text generation tasks Holtzman et al. (2020); Basu et al. (2021); Hewitt et al. (2022); Xu et al. (2023b). The evaluation of the methods in open-ended text generation tasks is future work.
We rely on automatic evaluations to evaluate the quality and the diversity of the generated texts. Human evaluation is desirable, especially for evaluating diversity. Although the automatic evaluation metrics for diversity used in this paper (e.g., P-BLEU, distinct-n, P-SentBERT) are shown to correlate with human evaluation, there is still a clear gap between automatic metrics and humans Tevet and Berant (2021).
DMBR and KMBR are much slower than DBS as they require the computation of the MBR objective which needs a computation of the utility function for $N^{2}$ times. Although recent work has shown that the computation of the MBR objective can be significantly reduced Cheng and Vlachos (2023); Jinnai and Ariu (2024); Deguchi et al. (2024); Vamvas and Sennrich (2024), it is not directly applicable to DMBR and KMBR. Reducing the inference time of these algorithms will be future work.
@@
We consider the Monte Carlo estimate (Eq. 7) as the target quality objective for simplicity. Exploring other quality objective functions such as model-based estimate Jinnai et al. (2024) is future work.
@@ Acknowledgements
We thank all the reviewers for their constructive comments throughout the manuscript review. We thank Naoto Ohsaka for providing insights on approximation algorithms for a non-monotonic submodular function maximization problem.
@@ Appendix A Proof of Submodularity
Appendix B Evaluation of the Coverage for CommonGen
The task encourages the generated sentence to contain as many input concepts as possible, so we also consider the coverage of the input concepts. The definition of coverage is the number of captured concepts divided by the number of input concepts. To compute the captured concepts, we use a Porter stemmer Porter (1980) to extract the stems of the input concepts and the generated sequences and compute the number of overlaps. Overall, the coverage is low ($<18\%$) for all the decoding algorithms. For $k=4$, the average coverage is $15.56,15.42,16.03,15.12,$ and $15.93$ for beam search, DBS-1.0, MBR, DMBR-1.0, and KMBR, respectively. We speculate this is because the text generation model (Zephyr-7b $\beta$) is not trained to follow the instructions on the lexical constraints without few-shot prompting.
Appendix C Examples of Generations
We show outputs generated by the decoding algorithms. Tables 2, 3, 4, 5, 6, and 7 are the examples of the generations with various sampling algorithms for each domain evaluated in Section 4. The examples show the generations of the first input source of each dataset.
@@ Appendix C Examples of Generations
Appendix D Evaluation of Oversampling Strategy
We additionally evaluate the performance of an oversampling strategy as a baseline. Oversampling strategy generates $N(>k)$ samples and then selects $k$ samples out of the $N$, maximizing the objective in Eq. (3):
Because $P_{\mathrm{human}}$ and $d_{\mathrm{human}}$ are inaccessible, we approximate them using a model and Eq. (10):
$H^{*}=\operatorname*{arg\,max}_{H\subseteq\mathcal{Y}}$ $\sum_{\mathbf{h}\in H}P(\mathbf{h}|\mathbf{x})-$
$\sum_{\mathbf{h}\in H}\sum_{\mathbf{h}^{\prime}\in H\setminus\{h\}}\frac{\lambda}{|H|}u(\mathbf{h},\mathbf{h}^{\prime}).$ (13)
The results with $N=128$ and $k=4$ are shown in Table 8. Overall, we observe that it performs slightly worse than DMBR, and the hyperparameter $\lambda$ is dependent on the generation probability of the sentences which in turn depends on sequence length (Table 9). The result indicates that the improvement of DMBR comes from the use of the utility function to select high-quality samples in addition to the oversampling.
Quality Diversity Decoder min BLEU $\uparrow$ mean BLEU $\uparrow$ max BLEU $\uparrow$ Pairwise-BLEU $\downarrow$ distinct-1 $\uparrow$ distinct-2 $\uparrow$ distinct-3 $\uparrow$ WMT’19 De-En ($k=4$) OS-0.1 30.96 35.09 39.44 75.92 0.28 0.33 0.34 OS-0.3 30.92 35.09 39.46 75.81 0.28 0.33 0.34 OS-0.5 30.92 35.09 39.46 75.74 0.28 0.33 0.34 OS-1.0 30.85 35.06 39.46 75.56 0.28 0.33 0.34 OS-2.0 30.83 35.05 39.50 75.37 0.28 0.33 0.34 MBR 31.25 35.17 39.20 75.14 0.28 0.33 0.35 DMBR-0.1 27.80 34.74 41.85 61.80 0.31 0.39 0.42 DMBR-0.3 21.21 32.03 43.51 40.74 0.37 0.50 0.55 DMBR-0.5 13.59 25.80 39.94 22.84 0.45 0.64 0.69 DMBR-1.0 10.76 20.59 33.41 15.54 0.51 0.71 0.75 DMBR-2.0 10.47 19.45 30.93 14.94 0.51 0.72 0.76 WMT’19 Ru-En ($k=4$) OS-0.1 27.59 31.59 35.62 75.68 0.27 0.33 0.34 OS-0.3 27.58 31.59 35.65 75.62 0.27 0.33 0.34 OS-0.5 27.58 31.59 35.65 75.62 0.27 0.33 0.34 OS-1.0 27.56 31.60 35.65 75.51 0.27 0.33 0.34 OS-2.0 27.50 31.57 35.66 75.38 0.27 0.33 0.34 MBR 28.08 32.28 36.44 75.12 0.27 0.33 0.35 DMBR-0.1 24.91 31.73 38.91 61.97 0.30 0.38 0.42 DMBR-0.3 19.74 29.60 40.33 41.30 0.35 0.50 0.55 DMBR-0.5 13.47 24.48 37.51 23.58 0.43 0.63 0.68 DMBR-1.0 10.49 19.84 31.51 16.23 0.48 0.70 0.75 DMBR-2.0 10.06 18.90 30.03 15.51 0.49 0.71 0.76
Table 8: Evaluation of the quality and diversity using the oversampling strategy (OS-$\lambda$) on WMT’19 De-En and Ru-En dataset (Appendix D). The size of the output $k$ is set to 4.
Appendix E Additional Figures and Tables
E.1 Additional Figures
The evaluation of the diversity as a function of min BLEU over the outputs on WMT’19 datasets is present in Figure 8. DMBR achieves a better trade-off than DBS and sampling algorithms with the same min BLEU score.
The Oracle (max) and the min quality scores on MS COCO, SQuADv2, CommonGen, and XSum with an output size of 4 are present in Figures 9 and 10. We observe similar trends as in machine translation tasks.
Figures 11 and 12 show the P-SentBERT as a function of the max and min BLEU and METEOR scores on MS COCO, SQuADv2, and CommonGen. The result indicates that DMBR and DBS are successfully generating diverse outputs, not only lexically but also semantically.
@@ E.1 Additional Figures
Figure 8: Min BLEU, P-BLEU, distinct-n on WMT’19 De-En and Ru-En. The size of the output $k$ is set to 4.
@@
Figure 10: Evaluation of the P-BLEU and distinct-2, 3 as a function of min BLEU (MS COCO), METEOR (SQuADv2, CommonGen), and ROUGE-L (XSum). The size of the output $k$ is set to 4.
@@
Figure 11: Evaluation of semantic textual similarity using P-SentBERT as a function of the Oracle quality score (max BLEU and max METEOR). The number of outputs $k$ is $4$.
(a) P-SentBERT $\downarrow$ (MS COCO)
(b) P-SentBERT $\downarrow$ (SQuADv2)
(c) P-SentBERT $\downarrow$ (CommonGen)
Figure 12: Evaluation of semantic textual similarity using P-SentBERT as a function of the Oracle quality score (min BLEU and min METEOR). The number of outputs $k$ is $4$.
E.2 Summary of the Results
Tables 9, 10, 11, 12, 13, 14, and 15 show the summary of the experimental results described in Section 4.